Night report 2022-11-16

Observers: Timo Kravtsov, Nada Ihanec

Support: Konrad Grzeciak, Mariusz Gromadzki

TAT: Michel Dennefeld, Lluís Galbany, Michael Fulton

EFOSC2 daytime Calibration

Evening: Bias, arcs+flats (gr 13, 11, 16, 18, 20)


Start: 00:15 UT, humidity 38%, wind speed 6.5 m/s, seeing 1.03" (DIMM)

Middle: 03:49 UT, humidity 28%, wind speed 3.5 m/s, seeing 0.78" (DIMM)

End: 08:20 UT, humidity 18%, wind speed 5.4 m/s, seeing 1.06" (DIMM)

There was a problem with a SOFI helium pump causing temperature and pressure to rise, but it was fixed during twilight. Observations started as planned with SOFI at the beginning of the night. At 1:15 UT SOFI cooling alarm was triggered again and the observations were halted. Switching to EFOSC at 2:20 UT, SOFI cooling problems persist. EFOSC focusing started at 2:30 UT. EFOSC refocusing started after standard star observations at 2:50 UT. EFOSC refocused at 8:00 UT. Observations ended at 08:36.

Night report 2022-11-16